# TorchCP: A Python Library for Conformal Prediction

### v4

@@ Abstract
Conformal prediction (CP) is a powerful statistical framework that generates prediction intervals or sets with guaranteed coverage probability. While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale DL scenarios. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning techniques, including DNN-based classifier/regressor, GNN, and LLM. Released under the LGPL-3.0 license, TorchCP comprises about 16k lines of code, validated with 100% unit test coverage and detailed documentation. Notably, TorchCP enables CP-specific training algorithms, online prediction, and GPU-accelerated batch processing, achieving up to 90% reduction in inference time on large datasets. With its low-coupling design, comprehensive suite of advanced methods, and full GPU scalability, TorchCP empowers researchers and practitioners to enhance uncertainty quantification across cutting-edge applications.
@@ 1 Introduction
Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty. While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated DL models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability for adapting to large-scale DL scenarios. This underscores the need for a comprehensive and scalable CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
This work presents TorchCP, a PyTorch-native library that advances conformal prediction for deep learning by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifier/regressor, GNN, and LLM. Its low-coupling design facilitates easy customization of components like trainers or scores, while full GPU support and batch processing ensure seamless scalability for large-scale tasks. TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix D.2), while also encompassing a broader algorithmic coverage than TorchUQ. This PyTorch-native integration and extensibility empower statisticians and DL practitioners to seamlessly apply CP across cutting-edge applications with enhanced DL compatibility.
@@ 2 Main contributions
Conformal prediction constructs prediction sets $C(X)$ for input $X$, ensuring the true outcome $Y$ satisfies $\mathbb{P}(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate. Existing CP libraries encounter significant challenges in deep learning, including limited native support for modern models (like DNN, GNN, and LLM), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications. TorchCP overcomes these shortcomings through a PyTorch-native framework that improves model compatibility, efficiency, and adaptability, thereby enhancing CP’s applicability to DL. Its low-coupling design and GPU-accelerated batch processing provide a robust foundation for tackling complex, real-world problems, detailed in the following parts.
@@ 2.1 Advantages for DL scenarios
Low-coupling design Low-coupling design is one of the most fundamental concepts in object-oriented design (McNatt and Bieman, 2001), which minimizes the dependency a class has on other classes. While many libraries (e.g., PUNCC) are implemented in a high-coupling design, TorchCP is structured into three core modules: Trainer, Score, and Predictor, each tailored to specific aspects of CP workflows. The Trainer module implements CP-specific training algorithms, such as ConfTr (Stutz et al., 2022), which optimizes deep models to improve efficiency. The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to training data. The Predictor module implements CP workflows, including calibration and prediction, to generate outputs, e.g., prediction sets/intervals, p-values, and conformal predictive distributions.
This low-coupling design is visualized in Figure 1, illustrating the interplay among modules. The Score module is used to create various score objects that serve as inputs for both the Trainer and Predictor modules. The Trainer module optimizes DL models using CP-specific loss functions. The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets, intervals, p-values, or predictive distributions. In this manner, TorchCP achieves strong reusability and extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s low-coupling design.
@@
DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
@@
Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifier/regressor, GNN, and LLM, along with their evaluation metrics in Table 4. Current libraries are typically limited in algorithmic coverage, e.g., MAPIE only includes split CP. In TorchCP, we provide 6 training algorithms, 17 scores, and 10 predictors. A brief overview of all supported methods and their advantages is available in Appendix A.
@@
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to deep learning applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)). Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
@@ 3 Related library
Conformal prediction is a distribution-free framework for quantifying uncertainty, yet existing libraries face notable limitations in deep learning applications. Libraries such as nonconformist, PUNCC, and MAPIE offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets. Moreover, tools like TorchUQ and Fortuna provide some GPU support but fall short in algorithmic coverage, e.g., TorchUQ supports only regressors. Additionally, many libraries like PUNCC are implemented in a high-coupling design, where predictors and score functions are tightly coupled within a single class. TorchCP addresses these gaps with its PyTorch-native integration, supporting a broader range of modern models and offering a modular, extensible architecture, as detailed in Table 8. We present a detailed comparison of the relevant libraries in Appendix F.
Table 2: Comparison of Open-Source CP Libraries  [^7] The definition of model support and functionality support is presented in Appendix F.. “Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests but is primarily designed for advanced deep learning models within the PyTorch ecosystem.. The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction. “$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
@@ Appendix A Details of Conformal Prediction Methods in TorchCP
Table 3: Advantages or use cases of CP methods across different tasks.
@@
Predictor Class-conditional CP To achieve coverage for every class
@@ A.1 Split CP
    Use the conformal threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
@@ A.2 DNN Classifier
C-Adapter (Liu et al., 2024): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss. The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
@@ A.3 DNN Regressor
NORABS (Papadopoulos et al., 2011; Papadopoulos and Haralambous, 2011): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\gamma+\xi(\bm{x})}$ normalizes the regression error using difficulty measure deviation $\xi(\bm{x})$. The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$. Additionally, we provide the normalized score for Gaussian Process Regression (Papadopoulos, 2024).
@@
Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their nonconformity scores using a specific aggregation function, such as mean or median. It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
@@ Appendix C Example Codes for TorchCP
TorchCP provides a collection of example codes in the “examples" folder  [^9] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks. In this section, we briefly introduce representative examples for the four supported tasks.
@@ C.2 Example codes for Semi-online conformal prediction
    DNN Classifier. For DNN classification tasks, this semi-online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added a semi-online conformal prediction example 151515 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
@@ D.1 Evaluation results
In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2). In Figure 3, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including Split CP, Class-conditional CP, and Clustered CP with APS. The number of trials is 1000.
@@
Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
@@ D.2.1 Classification Task
MNIST. We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing). In the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing. For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on a GPU. For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU. For Fortuna, we train the LeNet-5 model using the built-in training algorithm and perform model predictions with JAX on a GPU. We only report the total runtime for Fortuna as it ties the conformal calibration and inference processes together. Figure 6(a) presents the running times of the five libraries on the MINST dataset. The results show that TorchCP achieves the lowest calibration and inference time among the five libraries. Specifically, TorchCP completes calibration and inference for the entire dataset in just 0.07 seconds, using only 3.6% of the runtime needed by PUNCC (1.93 seconds). Moreover, the Crepes library exhibits the longest inference time on the MNIST dataset, likely due to its computation of p-values for each test sample’s non-conformity scores by comparing them against the entire calibration set. This substantial improvement highlights the efficiency advantage of TorchCP on small-scale datasets.
@@
Choices of Batch processing. To assess the efficiency of two processing modes (Full-batch and batch), we conduct experiments with TorchCP on both MNIST and ImageNet, using the same data splits and models as in prior experiments. We utilize GPU-accelerated computing for all experiments except the full-batch mode on ImageNet, where the dataset’s size exceeds GPU memory capacity. Figures 7(a) and  7(b) present the running times of different data processing modes on MNIST and ImageNet, respectively. On the small-scale dataset - MNIST, using full-batch processing on GPU achieves the fastest implementation among the three choices, requiring only 26% computational time of that with CPU. When it comes to the ImageNet dataset, GPU-based batch processing completes calibration and inference in approximately 104.29 seconds, compared to 1022.55 seconds for CPU-based processing, achieving a 9.8× speedup. These findings suggest that GPU-based full-batch processing is optimal for small-scale datasets where GPU memory is sufficient to load the whole dataset. Conversely, GPU-based batch processing is preferable for large-scale datasets like ImageNet, offering significant runtime advantages and enabling scalable, efficient deployment.
@@ D.2.2 Regression Task
We evaluate the computational efficiency and predictive performance of TorchCP and an R-based package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^16] https://cran.r-project.org/web/packages/opera/index.html. Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps. To assess computational efficiency, we apply the ACI algorithm(Gibbs and Candes, 2021) using both TorchCP and AdaptiveConformal. Both AdaptiveConformal and TorchCP are executed on CPU in this experiment to ensure a fair comparison. Moreover, since ACI performs recalibration for each test instance, the reported time includes both calibration and inference.

### v5

@@ Abstract
Conformal prediction (CP) is a powerful statistical framework that generates prediction intervals or sets with guaranteed coverage probability. While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale deep learning (DL) scenarios. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into DL techniques, including DNN-based classifiers/regressors, GNNs, and LLMs. Released under the LGPL-3.0 license, TorchCP comprises about 16k lines of code, validated with 100% unit test coverage and detailed documentation. Notably, TorchCP enables CP-specific training algorithms, online prediction, and GPU-accelerated batch processing, achieving up to 90% reduction in inference time on large datasets. With its low-coupling design, comprehensive suite of advanced methods, and full GPU scalability, TorchCP empowers researchers and practitioners to enhance uncertainty quantification across cutting-edge applications.
@@ 1 Introduction
Conformal prediction (CP) is a statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty. While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated deep learning (DL) models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability required for large-scale DL scenarios. This underscores the need for a comprehensive and scalable CP toolbox tailored for DL models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
This work presents TorchCP, a PyTorch-native library that advances conformal prediction for DL by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifiers/regressors, GNNs, and LLMs. Its low-coupling design facilitates easy customization of components like trainers or scores, while full GPU support and batch processing ensure seamless scalability for large-scale tasks. TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix F), while also encompassing a broader algorithmic coverage than TorchUQ. This PyTorch-native integration and extensibility empower statisticians and DL practitioners to seamlessly apply CP across cutting-edge applications with enhanced DL compatibility.
@@ 2 Main contributions
Conformal prediction constructs a prediction set $C(X)$ for an input $X$, ensuring the true outcome $Y$ satisfies $P(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate. Existing CP libraries encounter significant challenges in DL, including limited native support for modern models (like DNNs, GNNs, and LLMs), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications. TorchCP overcomes these shortcomings through a PyTorch-native framework that improves model compatibility, efficiency, and adaptability, thereby enhancing CP’s applicability to DL. Its low-coupling design and GPU-accelerated batch processing provide a robust foundation for tackling complex, real-world problems, detailed in the following parts.
@@ 2.1 Advantages for DL scenarios
Low-coupling design Low-coupling design is one of the most fundamental concepts in object-oriented design (McNatt and Bieman, 2001), which minimizes the dependency a class has on other classes. While many libraries (e.g., PUNCC) are implemented in a high-coupling design, TorchCP is structured into three core modules: Trainer, Score, and Predictor, each tailored to specific aspects of CP workflows. The Trainer module implements CP-specific training algorithms, such as ConfTr (Stutz et al., 2022), which optimizes deep models to improve efficiency. The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to the training data. The Predictor module implements CP workflows, including calibration and prediction, to generate outputs, e.g., prediction sets/intervals, p-values, and conformal predictive distributions.
This low-coupling design is visualized in Figure 1, illustrating the interplay among modules. The Score module is used to create various score objects that serve as inputs for both the Trainer and Predictor modules. The Trainer module optimizes DL models using CP-specific loss functions. The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets/intervals, p-values, or predictive distributions. In this manner, TorchCP achieves strong reusability and extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s low-coupling design.
@@
DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2025), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
@@
Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifiers/regressors, GNNs, and LLMs, along with evaluation metrics in Table 4. Current libraries are typically limited in algorithmic coverage, e.g., MAPIE only includes split CP. In TorchCP, we provide 6 training algorithms, 17 scores, and 10 predictors. A brief overview of all supported methods and their advantages is available in Appendix A.
@@
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to DL applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)). Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
@@ 3 Related library
Conformal prediction is a distribution-free framework for quantifying predictive uncertainty, yet existing libraries face notable limitations in DL applications. Libraries, such as nonconformist, PUNCC, and MAPIE, offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets. Moreover, tools like TorchUQ and Fortuna provide some GPU support but fall short in algorithmic coverage, e.g., TorchUQ supports only regressors. Additionally, many libraries like PUNCC are implemented in a high-coupling design, where predictors and score functions are tightly coupled within a single class. TorchCP addresses these gaps with its PyTorch-native integration, supporting a broader range of modern models and offering a modular, extensible architecture, as detailed in Table 8. We present a detailed comparison of the relevant libraries in Appendix F.
@@
Table 2: Comparison of Open-Source CP Libraries  [^7] The definitions of model support and functionality support are presented in Appendix F.. “Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests, but is primarily designed for advanced DL models within the PyTorch ecosystem.. The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction. “$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
@@ 4 Conclusion and Future Work
Acknowledgments and Disclosure of Funding
Hongxin Wei is supported by the Shenzhen Fundamental Research Program (Grant No. JCYJ20230807091809020). Bingyi Jing’s research is partly supported by NSFC 12371290. We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology for our research.
@@ Appendix A Details of Conformal Prediction Methods in TorchCP
Predictor Class-conditional CP To achieve coverage for each class
@@
Table 3: Advantages or use cases of CP methods across different tasks.
@@ A.1 Split CP
    Use the threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
@@ A.2 DNN Classifier
C-Adapter (Liu et al., 2025): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss. The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
@@ A.3 DNN Regressor
NORABS (Papadopoulos et al., 2011; Papadopoulos and Haralambous, 2011): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\gamma+\xi(\bm{x})}$ normalizes the regression error using difficulty measure deviation $\xi(\bm{x})$. The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$. Additionally, we provide the normalized score for Gaussian Process Regression (Papadopoulos, 2024).
@@
Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their non-conformity scores using a specific aggregation function, such as mean or median. It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
@@ Appendix C Example Codes for TorchCP
TorchCP provides a collection of example codes in the “examples” folder  [^9] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks. In this section, we briefly introduce representative examples for the four supported tasks.
@@ C.2 Example codes for semi-online conformal prediction
    DNN Classifier. For DNN classification tasks, this semi-online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added a semi-online conformal prediction example  [^15] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
@@ D.1 Evaluation results
In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS ($\lambda$=0.1, $k$ =0), and SAPS ($\lambda$=0.2). Figure 3 compares Split CP, Class-conditional CP, and Clustered CP on the ImageNet dataset, each employing the APS scoring rule. Results are averaged over 1,000 independent trials.
@@
Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
@@ D.2.1 Classification Task
MNIST. We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing). For the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing. For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on the GPU. For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU. For Fortuna, we train the LeNet-5 model using the built-in training algorithm and perform model predictions with JAX on a GPU. We only report the total runtime for Fortuna as it ties the conformal calibration and inference processes together. Figure 6(a) presents the running times of the five libraries on the MINST dataset. The results show that TorchCP achieves the lowest calibration and inference time among the five libraries. Specifically, TorchCP completes calibration and inference for the entire dataset in just 0.07 seconds, using only 3.6% of the runtime needed by PUNCC (1.93 seconds). Moreover, the Crepes library exhibits the longest inference time on the MNIST dataset, likely due to its computation of p-values for each test sample’s non-conformity scores by comparing them against the entire calibration set. This substantial improvement highlights the efficiency advantage of TorchCP on small-scale datasets.
@@
[…]
@@ D.2.2 Regression Task
We evaluate the computational efficiency and predictive performance of TorchCP and an R-based CP package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^16] https://cran.r-project.org/web/packages/opera/index.html. Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps. To assess computational efficiency, we apply the ACI algorithm(Gibbs and Candes, 2021) using both TorchCP and AdaptiveConformal. Both AdaptiveConformal and TorchCP are executed on CPU in this experiment to ensure a fair comparison. Moreover, since ACI performs recalibration for each test instance, the reported time includes both calibration and inference.